????
The Travel Time bus company has determined that when x tourists are given a particular bus
tour, the profit P is given by
P(x) = –0.25 x2 + 29.50x – 170
(a) What is the company’s profit if 42 tourists are given the tour?
(b) How many tourists should be given the tour in order to maximize the company’s profit?

Answer by nerdybill(7384)

The Travel Time bus company has determined that when x tourists are given a particular bus
tour, the profit P is given by
P(x) = –0.25 x^2 + 29.50x – 170
(a) What is the company’s profit if 42 tourists are given the tour?
set x = 42 and solve for P(42):
P(42) = –0.25 (42)^2 + 29.50(42) – 170
P(42) = –441 + 1239 – 170
P(42) = 628
.
(b) How many tourists should be given the tour in order to maximize the company’s profit?
Max is at the vertex:
x = -b/(2a)
x = -29.50/(2(–0.25))
x = -29.50/(-0.5)
x = 59 tourists